# Short star products for quantum symmetric pairs and applications

### 06 Mar 2026 v1

@@ 1.1. Short star products
The concept of short star products for commutative graded Poisson algebras originated in the context of conformal field theory in the work of Beem, Peelaers and Rastelli [BPR17] and was further developed mathematically by Etingof and Stryker [ES20]. Any $\mathbb{N}$}-graded algebra $A=\bigoplus_{k\in{\mathbb{N}}}A_{n}$ is filtered with ${\mathcal{F}}_{n}(A)=\bigoplus_{k=0}^{n}A_{k}$. A filtered deformation of $A$ is an additional associative algebra structure $\ast:A\otimes A\rightarrow A$ which turns $(A,{\mathcal{F}},\ast)$ into a filtered algebra such that the associated graded algebra ${\mathrm{gr}}_{\mathcal{F}}(A,\ast)$ is isomorphic to $A$ under the canonical isomorphism of vector spaces ${\mathrm{gr}}_{\mathcal{F}}(A)\cong A$. In this situation, the new algebra structure $\ast$ is called a star product on $A$. By associativity, any filtered deformation of $A$ gives rise to a Poisson bracket on the commutative algebra $A$.
The papers [BPR17] and [ES20] identify a crucial concept for filtered deformations: A star product $\ast$ on $\mathcal{A}$} is called short if
@@
The essence of shortness is the nonzero lower bound in the direct sum in (1.1) which poses a strong restriction on the possible filtered deformations. In [ES20] short star products on commutative, graded Poisson algebras were identified as a new, important structure in representation theory.
Filtered deformations can be obtained via quantization maps. Let $(B,{\mathcal{G}})$ be a filtered associative algebra. A quantization map is an isomorphism of filtered vector spaces $\phi:A\rightarrow B$ such that the associated graded map ${\mathrm{gr}}(\phi):{\mathrm{gr}}_{\mathcal{F}}(A)\cong A\rightarrow{\mathrm{gr}}_{\mathcal{G}}(B)$ is an isomorphism of graded algebras. If $\phi$ is a quantization map then
@@ 1.4. Results
Let ${\mathcal{W}}_{X,\tau}$ be the quotient of $U^{\mathrm{poly}}$ by the two sided ideal generated by all $K_{i}^{-1}$ for $i\in I\setminus X$. We call ${\mathcal{W}}_{X,\tau}$ the parabolic Heisenberg double corresponding to the (generalized) Satake diagram $(X,\tau)$. We obtain the commutative diagram
@@
The projection $\eta$ turns ${\mathcal{A}}_{X,\tau}^{-}$ into a left ${\mathcal{W}}_{X,\tau}$-module which can be thought of as a Fock space for the parabolic Heisenberg double. The algebra ${\mathcal{W}}_{X,\tau}$ has a triangular decomposition
@@ 1.4.4. Application 4: The quasi $K$-matrix for ${\mathcal{B}}_{\mathbf{c}}$
Coideal property of ${\mathcal{B}}_{\mathbf{c}}$ Homogeneity of the image of ${\mathcal{B}}_{\mathbf{c}}$ in the parabolic Heisenberg double ${\mathcal{W}}_{X,\tau}$ ${\mathcal{B}}_{\mathbf{c}}\cong({\mathcal{A}}^{-}_{X,\tau},*)$ for a ${\mathbb{Z}}/2$-equivariant short star product Conversion of left to right *-actions of $({\mathcal{A}}^{-}_{X,\tau})_{1}$ on ${\mathcal{A}}^{-}_{X,\tau}$ $\sigma_{\tau}:{\mathcal{B}}_{\mathbf{c}}\rightarrow{\mathcal{B}}_{\mathbf{c}}$ is an anti-automorphism $\sigma\circ\tau$ fixes $\partial^{L}_{i}(T_{w_{X}}^{-1}(E_{i}))$ and $\partial^{R}_{i}(T_{w_{X}}(E_{i}))$ for $i\notin X$ and their $F$-versions Construction of the bar involution on ${\mathcal{B}}_{\mathbf{c}}$ $(\psi^{-1}\otimes\mathrm{id})(\Theta)$ is the tensor quasi $K$-matrix for $({U_{q}(\mathfrak{g})},{\mathcal{B}}_{\mathbf{c}})$
@@ 1.5. Organization
The main body of this paper is organized in four sections. In Section 2 we fix notation for quantized enveloping algebras and quantum symmetric pairs. We spell out the properties of various versions ${\mathcal{R}}_{X}^{\pm,r}$, ${\mathcal{R}}_{X}^{\pm,l}$ of quantum parabolic nilradicals which are obtained as coinvariants for suitable coactions on quantum standard parabolics. The interplay between the different versions of quantum parabolic nilradicals will be crucial in later sections. We discuss the relation between QSP-subalgebras and quantum horospherical subalgebras and revisit the Letzter map from [KY21]. We then define the parabolic Heisenberg double ${\mathcal{W}}_{X,\tau}$ and prove Theorem A as a consequence of the coideal property of the QSP-subalgebra ${\mathcal{B}}_{\mathbf{c}}$.
@@ 2. The Letzter map
We introduce background and notation for quantum groups $U_{q}(\mathfrak{g})$} and quantum symmetric pairs $({U_{q}(\mathfrak{g})},{\mathcal{B}}_{\mathbf{c}})$ and recall the Letzter map. We will see that the Letzter map factors over the parabolic Heisenberg double ${\mathcal{W}}_{X,\tau}$. The algebra ${\mathcal{W}}_{X,\tau}$ has a natural vector space grading. As a first main result we will show in Theorem 2.14 that the image of ${\mathcal{B}}_{\mathbf{c}}$ in ${\mathcal{W}}_{X,\tau}$ is a graded subspace. This fact will be a crucial ingredient when proving the shortness of the star product in Section 3.3
@@ 2.5. The Letzter map
2.6. The parabolic Heisenberg double
@@
the parabolic Heisenberg double associated to the generalized Satake diagram $(X,\tau)$. Equation (2.41) implies that
@@ 3.1. Star products
    $a*b\in\bigoplus_{k=0}^{\lfloor\frac{m+n}{2}\rfloor}A_{m+n-2k}\qquad\mbox{for all $a\in A_{m},b\;A_{n}$.}$
@@
    The linear map $C_{1}:A\otimes A\rightarrow A$ of degree 2-2 defines a Hochschild $2$-cocycle, that is, $aC_{1}(b,c)-C_{1}(ab,c)+C_{1}(a,bc)-C_{1}(a,b)c=0$ for all $a,b,c\in A$.
@@
    If the graded algebra $A$ is a commutative Poisson algebra with Poisson bracket $\{\,,\,\}:A\otimes A\rightarrow A$ of degree 2-2 and $\ast:A\times A\rightarrow A$ is a ${\mathbb{Z}}/2$-equivariant $\ast$-product on $A$ such that
    (3.2) $\{a,b\}=C_{1}(a,b)-C_{1}(b,a)\qquad\mbox{for all $a,b\in A$}$
[…]
@@
    Nondegenerate short star products on connected graded algebras are classified in terms nondegenerate twisted traces [ES20, Proposition 3.3]. In this paper we are concerned with the case where $A$ is noncommutative and $\dim A_{0}>1$.
@@ 4.4. The bar involution for ${\mathcal{B}}_{\mathbf{c}}$
(4.18) $\sigma\circ T^{\prime\prime}_{i,e}$ $=T^{\prime}_{i,e}\circ\sigma,$
@@
$=(-1)^{k}q^{\frac{1}{2}((\alpha,\alpha)-\sum_{n=1}^{k}(\alpha_{j_{n}},\alpha_{j_{n}}))}K_{-\tau(\alpha)}\omega\circ\tau(f),$
@@
$=(-1)^{k}q^{\frac{1}{2}((\alpha,\alpha)-\sum_{n=1}^{k}(\alpha_{j_{n}},\alpha_{j_{n}}))}\omega\circ\tau(f)K_{\tau(\alpha)}.$

### 25 Mar 2026 v2

@@ 1.1. Short star products
The concept of short star products for commutative graded Poisson algebras originated in the context of conformal field theory in the work of Beem, Peelaers and Rastelli [BPR17] and was further developed mathematically by Etingof and Stryker [ES20]. Any $\mathbb{N}$}-graded algebra $A=\bigoplus_{k\in{\mathbb{N}}}A_{n}$ is filtered with ${\mathcal{F}}_{n}(A)=\bigoplus_{k=0}^{n}A_{k}$. A star product on $A$ is an additional associative algebra structure $\ast:A\otimes A\rightarrow A$ which turns $(A,{\mathcal{F}},\ast)$ into a filtered algebra such that the associated graded algebra ${\mathrm{gr}}_{\mathcal{F}}(A,\ast)$ is isomorphic to $A$ under the canonical isomorphism of vector spaces ${\mathrm{gr}}_{\mathcal{F}}(A)\cong A$. Recall that if $(B,{\mathcal{G}})$ is a filtered algebra then the associated graded algebra ${\mathrm{gr}}_{\mathcal{G}}(B)$, if commutative, naturally carries a Poisson structure. In this case $(B,{\mathcal{G}})$ is called a filtered quantization of the commutative graded Poisson algebra ${\mathrm{gr}}_{\mathcal{G}}(B)$. Star products on commutative graded algebras $A$ are hence particular examples of filtered quantizations of $A$.
The papers [BPR17] and [ES20] identify a crucial concept for star products: A star product $\ast$ on $A$ is called short if
@@
The nonzero lower bound in the direct sum in (1.1) poses a strong restriction on the possible filtered deformations. In [ES20] short star products on commutative, graded Poisson algebras were identified as a new, important structure in representation theory.
Star products on $A$ can be obtained via quantization maps. Let $(B,{\mathcal{G}})$ be a filtered associative algebra. A quantization map is an isomorphism of filtered vector spaces $\phi:A\rightarrow B$ such that the associated graded map ${\mathrm{gr}}(\phi):{\mathrm{gr}}_{\mathcal{F}}(A)\cong A\rightarrow{\mathrm{gr}}_{\mathcal{G}}(B)$ is an isomorphism of graded algebras. If $\phi$ is a quantization map then
@@ 1.4. Results
Let ${\mathcal{W}}_{X,\tau}$ be the quotient of $U^{\mathrm{poly}}$ by the two sided ideal generated by all $K_{i}^{-1}$ for $i\in I\setminus X$. We call ${\mathcal{W}}_{X,\tau}$ the horospherical Heisenberg double corresponding to the (generalized) Satake diagram $(X,\tau)$. We obtain the commutative diagram
@@
The projection $\eta$ turns ${\mathcal{A}}_{X,\tau}^{-}$ into a left ${\mathcal{W}}_{X,\tau}$-module which can be thought of as a Fock space for the horospherical Heisenberg double. The algebra ${\mathcal{W}}_{X,\tau}$ has a triangular decomposition
@@ 1.4.4. Application 4: The quasi $K$-matrix for ${\mathcal{B}}_{\mathbf{c}}$
Coideal property of ${\mathcal{B}}_{\mathbf{c}}$ Homogeneity of the image of ${\mathcal{B}}_{\mathbf{c}}$ in the horospherical Heisenberg double ${\mathcal{W}}_{X,\tau}$ ${\mathcal{B}}_{\mathbf{c}}\cong({\mathcal{A}}^{-}_{X,\tau},*)$ for a ${\mathbb{Z}}/2$-equivariant short star product Conversion of left to right *-actions of $({\mathcal{A}}^{-}_{X,\tau})_{1}$ on ${\mathcal{A}}^{-}_{X,\tau}$ $\sigma_{\tau}:{\mathcal{B}}_{\mathbf{c}}\rightarrow{\mathcal{B}}_{\mathbf{c}}$ is an anti-automorphism $\sigma\circ\tau$ fixes $\partial^{L}_{i}(T_{w_{X}}^{-1}(E_{i}))$ and $\partial^{R}_{i}(T_{w_{X}}(E_{i}))$ for $i\notin X$ and their $F$-versions Construction of the bar involution on ${\mathcal{B}}_{\mathbf{c}}$ $(\psi^{-1}\otimes\mathrm{id})(\Theta)$ is the tensor quasi $K$-matrix for $({U_{q}(\mathfrak{g})},{\mathcal{B}}_{\mathbf{c}})$
@@ 1.5. Organization
The main body of this paper is organized in four sections. In Section 2 we fix notation for quantized enveloping algebras and quantum symmetric pairs. We spell out the properties of various versions ${\mathcal{R}}_{X}^{\pm,r}$, ${\mathcal{R}}_{X}^{\pm,l}$ of quantum parabolic nilradicals which are obtained as coinvariants for suitable coactions on quantum standard parabolics. The interplay between the different versions of quantum parabolic nilradicals will be crucial in later sections. We discuss the relation between QSP-subalgebras and quantum horospherical subalgebras and revisit the Letzter map from [KY21]. We then define the horospherical Heisenberg double ${\mathcal{W}}_{X,\tau}$ and prove Theorem A as a consequence of the coideal property of the QSP-subalgebra ${\mathcal{B}}_{\mathbf{c}}$.
@@ 2. The Letzter map
We introduce background and notation for quantum groups $U_{q}(\mathfrak{g})$} and quantum symmetric pairs $({U_{q}(\mathfrak{g})},{\mathcal{B}}_{\mathbf{c}})$ and recall the Letzter map. We will see that the Letzter map factors over the horospherical Heisenberg double ${\mathcal{W}}_{X,\tau}$. The algebra ${\mathcal{W}}_{X,\tau}$ has a natural vector space grading. As a first main result we will show in Theorem 2.14 that the image of ${\mathcal{B}}_{\mathbf{c}}$ in ${\mathcal{W}}_{X,\tau}$ is a graded subspace. This fact will be a crucial ingredient when proving the shortness of the star product in Section 3.3
@@ 2.5. The Letzter map
2.6. The horospherical Heisenberg double
@@
the horospherical Heisenberg double associated to the generalized Satake diagram $(X,\tau)$. Equation (2.41) implies that
@@ 3.1. Star products
    (3.2) $a*b\in\bigoplus_{k=0}^{\lfloor\frac{m+n}{2}\rfloor}A_{m+n-2k}\qquad\mbox{for all $a\in A_{m},b\in A_{n}$.}$
@@
    The linear map $C_{1}:A\otimes A\rightarrow A$ of degree 2-2 defines a Hochschild $2$-cocycle, that is, $aC_{1}(b,c)-C_{1}(ab,c)+C_{1}(a,bc)-C_{1}(a,b)c=0$ for all $a,b,c\in A$. If the graded algebra $A$ is commutative, then
    $\{a,b\}:=C_{1}(a,b)-C_{1}(b,a)\qquad\mbox{for all $a,b\in A$}$
    defines a Poisson structure on $A$ and we think of the algebra (A,)(A,*) as of a quantization of the Poisson algebra $(A,\{.,.\})$.
@@
    In [ES20] all star products are assumed to satisfy the condition (3.2) and the term ${\mathbb{Z}}/2$-equivariant is used in relation to filtered quantizations equipped with quantization maps, see [ES20, Proposition 2.5].
@@
    Nondegenerate short star products on connected graded algebras are classified in terms nondegenerate twisted traces on filtered quantizations [ES20, Proposition 3.3]. In this paper we are concerned with the case where $A$ is noncommutative and $\dim A_{0}>1$.
@@ 4.4. The bar involution for ${\mathcal{B}}_{\mathbf{c}}$
(4.18) $\sigma\circ T^{\prime\prime}_{i,e}$ $=T^{\prime}_{i,-e}\circ\sigma,$
@@
$=q^{\frac{1}{2}((\alpha,\alpha)-\sum_{n=1}^{k}(\alpha_{j_{n}},\alpha_{j_{n}}))}K_{-\tau(\alpha)}\omega\circ\tau(f),$
@@
$=q^{\frac{1}{2}((\alpha,\alpha)-\sum_{n=1}^{k}(\alpha_{j_{n}},\alpha_{j_{n}}))}\omega\circ\tau(f)K_{\tau(\alpha)}.$